OpenStudy (anonymous):

What is the relationship of a system of equations which have the same slope but different y intercepts?

OpenStudy (cwrw238):

they are parallel to each other

OpenStudy (cwrw238):



OpenStudy (cwrw238):

assuming the bottom equation is y = x then next one will be y = x + 1 top one y = x + 2 the 3 all have the same slope (1) and are therefore parallel